OpenStudy (anonymous):

graph the equation using the slope and the y-intercept y=5/3x+3

OpenStudy (dumbcow):

y=mx+b m=slope b=y-intercept First plot a point for y_intercept (0,b) Then use slope to get second point (rise/run)

OpenStudy (anonymous):

would it be y=3

OpenStudy (dumbcow):

yes good

OpenStudy (anonymous):

(0,3) (-3,0)

OpenStudy (anonymous):

would that be the answer

OpenStudy (dumbcow):

(0,3) is correct but not your second point lets look at slope 5/3 right?

OpenStudy (dumbcow):

slope = rise/run = 5/3 from any point if you go up 5 and to right 3 you will be on correct line

OpenStudy (anonymous):

Or you can think about plugging in 0 for y you have \[0 = \frac{5}{3} x + 3\] So solve for x.